40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 269     Entry time: Tue Aug 10 21:40:42 2010
Author: Megan and Frank 
Type: DailyProgress 
Category:  
Subject: Both Cavities Locked! 

We got both cavities locked today!

They stayed locked for at least an hour and were still locked when we left. Hopefully they'll still be locked tomorrow morning!

We aligned both beams into the photodiode (they were already pre-aligned) and maximized the alignment to get the largest signal out of the photodiode. We locked the signal from the photodiode with the IFR/Marconi 2023B signal generator and took some measurements of noise and transfer functions to determine UGFs, which I will post graphs of when I put them together. The plan is to let the temperatures really settle down (they were still settling a bit when we left) and take a good measurement of the noise and measure the transfer functions to determine UGF tomorrow. Then we will pull everything apart (not really) to replace the two cables that have extensions on them to see if something in the connection can be fixed to give better stabilization.

*I added a graph of the noise we measured yesterday at 3 different times*

Attachment 1: BeatSignal.png  17 kB  Uploaded Wed Aug 11 12:24:35 2010  | Hide | Hide all
BeatSignal.png
ELOG V3.1.3-